# Efficient Causal Graph Discovery Using Large Language Models

### v3

@@ Abstract
LLM, large language models, causality, machine learning, ICML
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved AI capabilities (OpenAI, 2023). LLMs have access to a vast amount of training data, allowing them to acquire a substantial amount of knowledge, ranging from common-sense to specific domains, such as math and science. Moreover, recent papers have argued that complex behaviors, such as writing code, generating long stories, and even some reasoning capabilities, can emerge from such large-scale training (Wei et al., 2022; Rozière et al., 2023; Zhao et al., 2023; Yao et al., 2023). Impressively, LLMs can use their knowledge to generate plausible explanations given observations and even reason on counterfactuals (Wang et al., 2023; Li et al., 2023; Kıcıman et al., 2023). Given these emergent abilities to understand and learn relationships from textual corpora, a natural question arises: Can LLMs reason about causal relationships?
@@
Causal discovery involves identifying the causal relationships between different variables (Pearl, 2009; Peters et al., 2017). This process usually results in a Directed Acyclic Graph (DAG), with edges representing the existence and direction of causal influences. This causal graph represents the relationships between variables which govern the data generation process. The causal graph also allows us to determine how changes in one variable might influence others, and serves as the foundation for further analysis related to specific tasks, including effect inference, prediction, or attribution. Ensuring the correctness of this causal graph is essential for the reliability of any subsequent analysis.
@@
Traditional causal discovery frameworks (Bello et al., 2023; Zheng et al., 2018; Meek, 2023) rely on numerical observations and use statistical or continuous optimization based methods to unravel confounding and causal factors. While such rigorous numerical and abstract reasoning abilities remain challenges for LLMs (Gendron et al., 2023; Xu et al., 2023; Imani et al., 2023), recent works have shown that LLMs are capable of leveraging their internal knowledge to perform commonsense and simple logical as well as causal reasoning (Wei et al., 2022; Zhao et al., 2023; Yao et al., 2023; Kıcıman et al., 2023). Inspired by these results, the proposed framework utilizes LLMs to respond to causal queries solely based on metadata, such as variable names and descriptions, without accessing the numerical observations. This is similar to the metadata and knowledge-based reasoning that is done by human domain experts when they construct causal graphs.
@@
Previous works (Kıcıman et al., 2023; Choi et al., 2022; Long et al., 2023b) have shown that LLMs are capable of inferring pairwise causal relationships given only metadata. However, extending this pairwise framework to full graph discovery presents additional challenges. One straightforward approach which previous works use is to perform pairwise queries over all possible pairs of variables. However, this method is not scalable since it requires ($n\choose 2$ total queries where $n$ is the number of variables. Thus, the number of queries required grows quadratically with the number of causal variables. As a result, the experiments in these works are limited to very small causal graphs.
@@
To the best of our knowledge, we present the first LLM-based causal graph discovery method applicable to larger graphs. The proposed method is much more efficient than previous LLM-based methods, and does not require observational data (unlike numerical methods), while still achieving state-of-the-art or competitive performance on three graphs of varying sizes. In addition, we show how observational data can easily be incorporated when available to improve the proposed method’s performance. We provide code to reproduce all experiments at this URL.
@@ 2 Related Work
Causal graph discovery has been extensively studied in many settings (Zanga & Stella, 2023). Until recently, most proposed methods were statistical methods that used observational data to infer causal relationships between variables (Bello et al., 2023; Zheng et al., 2018; Meek, 2023; Spirtes et al., 1993). Recently, many works have proposed to use LLMs for causal graph discovery (Kıcıman et al., 2023; Choi et al., 2022; Long et al., 2023b). These LLM-based methods do not rely on observational data. Instead, they utilize metadata associated with the input variables, specifically variable descriptions in natural language. This approach closely resembles how human experts construct causal graphs using their domain knowledge. In the rest of this section, we provide a brief overview of popular causal graph discovery algorithms (both statistical and LLM-based).
2.1 Statistical Methods
2.1.1 Peter-Clark Algorithm (PC) (Spirtes & Glymour, 1991)
PC is a constraint-based algorithm which starts by defining a complete undirected graph over the variables. It then performs a series of conditional independence tests and uses these tests to remove the associated edges from the complete graph. Finally, it orients the edges remaining in the graph using rules that take into account acyclicity and other causal assumptions.
2.1.2 Greedy Equivalence Search (GES) (Meek, 2023)
GES is a score-based causal discovery algorithm, meaning that it seeks to optimize some score that is a function of the graph. In our case, the score is the Bayesian Information Criterion (BIC) (Schwarz, 1978). GES begins with an empty graph. It then alternates between 2 phases: the forward phase and the backward phase. During the forward phase, the edge that, once added, makes the graph’s score go up the most is repeatedly added until there is no edge that increases the score. During the backward phase, the edge that, once removed, makes the graph’s score go up the most, is repeatedly removed until there is no edge that increases the score. The algorithm finishes once a local maximum is reached during the backward phase.
2.1.3 NOTEARS (Zheng et al., 2018)
Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) assumes that the data is generated linearly according to the model $\textbf{X}=\textbf{X}W+N$ where $N$ is a noise vector with independent entries, X is a $n\times d$ observation matrix containing $n$ i.i.d observations of $d$ variables and $W_{ij}$ can be viewed as the probability that variable $V_{i}$ has a direct causal influence on $V_{j}$, or equivalently, that there is a directed edge from node $i$ to node $j$ in the causal graph between the $d$ variables.
NOTEARS thus solves the following constrained optimization problem using the augmented Lagrangian method:
$\hat{W}=argmin_{W}\frac{1}{2n}||\textbf{X}-\textbf{X}W||^{2}_{F}+\lambda||W||_{1}$
s.t.
$exp(W\circ W)-d=0$
where $\lambda||W||_{1}$ is a regularizer and the differentiable constraint $exp(W\circ W)-d=0$ is added to ensure that the resulting causal graph is a directed acyclic graph (DAG) as desired.
The entries $W_{ij}$ in $W$ which are above a certain threshold are then added to the output causal graph as directed edges from node $i$ to node $j$.
2.1.4 DAGMA (Bello et al., 2023)
Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) solves the same optimization problem as NOTEARS but using a different acyclicity constraint based on the log-determinant function.
2.2 LLM-based methods
2.2.1 Pairwise Method
Kıcıman et al. (2023); Choi et al. (2022); Long et al. (2023b) use pairwise queries to infer the causal relationship between 2 variables at a time. Given variables $A$ and $B$, there are three possible causal relationships, $A\rightarrow B$, $A\leftarrow B$, and no relationship. This query can be formulated as a question in natural language with the following prompt:
@@ 2.2.1 Pairwise Method
Given the above information, which of
the following is the most likely:
@@ B. Changing <B> causes a change in <A>
C. There is no causal relationship
between <A> and <B>
@@ 3.1 Main Method
You are a helpful assistant to a
neuropathic pain diagnosis expert.
The following factors are key
variables related to neuropathic pain
diagnosis which have various causal
effects on each other. Our goal is
to construct a causal graph between
these variables.
@@
Now you are going to use the data to
construct a causal graph. You will
start with identifying the variable(s)
that are unaffected by any other
variables.
Think step by step. Then, provide your
final answer (variable names only)
within the tags <Answer>...</Answer>.
@@
Given <Independent Variables> is(are)
not affected by any other variable and
the following causal relationships:
@@
Select the variables that are caused by
<Currently Visited Node>.
Think step by step. Then, provide your
final answer (variable names only)
within the tags <Answer>...</Answer>.
@@ 3.2 Adding Statistics to the Prompt
Additionally, the Pearson Correlation
Coefficients between <Currently Visited
Node> and the other variables are as
follows:
<A>: <Pearson Correlation Coefficient
     between <Currently Visited Node>
     and A>
<B>: <Pearson Correlation Coefficient
     between <Currently Visited Node>
     and B>
@@
Additionally, the Pearson Correlation
Coefficient between <A> and <B> is
<Correlation Coefficient>.
@@ 4.1 Experimental Setup
We use the GPT-4 API (OpenAI, 2023) for all of our experiments. We evaluate the proposed approach on three causal graphs of varying sizes:
@@
    Asia (Lauritzen & Spiegelhalter, 1988). The Asia causal graph is a small causal graph that models lung cancer in patients having potentially recently visited Asia. The ground-truth graph has 8 nodes and 8 edges. Some examples of variables in the graph are Visit to Asia, Smoking, Bronchitis, and Lung Cancer.
@@
    Child (Spiegelhalter et al., 1993). The Child causal graph is a medium-sized causal graph which models congenital heart disease in newborn babies. The ground-truth graph has 20 nodes and 25 edges. Some examples of variables in the graph are Birth Asphyxia, Lung Flow, and Chest XRay.
@@
    Neuropathic Pain (Tu et al., 2019) The Neuropathic Pain causal graph is a very large causal graph consisting of 221 nodes and 770 edges. It contains relationships between different nerves and the associated symptoms that patients express. It mainly consists of symptom diagnoses, pattern diagnoses, and pathophysiological diagnoses. The symptom diagnoses describe the discomfort of patients. The pattern diagnoses identify nerve roots that lead to symptom patterns. The pathophysiological diagnoses identify discoligamentous injuries.
@@
We get the Asia and Child causal graphs and data from the bnlearn package (Scutari, 2010), and their descriptions from Long et al. (2023a). We get the Neuropathic Pain causal graph data and descriptions from Tu et al. (2019).
We compare the proposed method, as well as the proposed method plus observational statistics from 1000 and 10000 samples, against the numerical methods of GES, PC, NOTEARS, and DAGMA (all discussed in section 2.2), and against pairwise queries, and pairwise queries plus observational statistics from 1000 and 10000 samples. We use the implementations of GES and PC from the causal-learn package (Zheng et al., 2023), the implementation of NOTEARS from Zheng et al. (2018), and the implementation of DAGMA from Bello et al. (2023). For the numerical methods, we conduct experiments in the low (100 samples), medium (1000 samples), and high (10000 samples) data regimes.
@@ 4.2 Metrics
Method Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F Score $\uparrow$ NPE NHD $\downarrow$ Ref. NHD NHD Ratio $\downarrow$
Numerical Methods
100 Samples
GES 0.23 0.38 0.38 0.38 8 0.28 0.44 0.63
PC 0.33 0.75 0.38 0.5 4 0.24 0.48 0.5
NOTEARS ($\lambda=0.02$) 0.5 1 0.5 0.67 4 0.11 0.33 0.33
DAGMA ($\lambda=0.02$) 0.22 0.67 0.25 0.36 3 0.28 0.44 0.64
1000 Samples
GES 0.67 0.86 0.75 0.8 7 0.06 0.31 0.2
PC 0.5 0.71 0.63 0.67 7 0.1 0.31 0.33
NOTEARS ($\lambda=0.005$) 0.44 0.8 0.5 0.62 5 0.10 0.27 0.38
DAGMA ($\lambda=0.01$) 0.5 1 0.5 0.67 4 0.11 0.33 0.33
10000 Samples
GES 0.7 0.78 0.88 0.82 9 0.06 0.35 0.18
PC 0.55 0.67 0.75 0.71 9 0.1 0.35 0.29
NOTEARS ($\lambda=0.005$) 0.44 0.8 0.5 0.62 5 0.14 0.36 0.38
DAGMA ($\lambda=0.005$) 0.5 1 0.5 0.67 4 0.11 0.33 0.33
LLM Methods
Pairwise 0.33 0.35 0.88 0.5 20 0.22 0.44 0.5
Pairwise + 1000 samples 0.38 0.43 0.75 0.54 14 0.20 0.45 0.45
Pairwise + 10000 samples 0.47 0.47 1 0.64 17 0.14 0.39 0.36
Ours 0.88 1 0.88 0.93 7 0.016 0.23 0.067
Ours + 1000 samples 0.8 0.8 1 0.89 10 0.031 0.28 0.11
Ours + 10000 samples 0.8 0.8 1 0.89 10 0.031 0.28 0.11
Table 1: Results on the Asia causal graph (8 nodes, 8 edges). The proposed method with no observational statistics outperforms all other methods with a F score of 0.93 and a NHD ratio of 0.067.
Method Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F Score $\uparrow$ NPE NHD $\downarrow$ Ref. NHD NHD Ratio $\downarrow$
Numerical Methods
100 Samples
GES 0.24 0.41 0.36 0.38 22 0.11 0.18 0.62
PC 0.19 0.37 0.28 0.32 19 0.12 0.17 0.68
NOTEARS ($\lambda=0.01$) 0.14 0.26 0.24 0.25 23 0.1 0.13 0.75
DAGMA ($\lambda=0.005$) 0.14 0.2 0.32 0.24 41 0.13 0.17 0.76
1000 Samples
GES 0.31 0.41 0.47 0.44 34 0.08 0.15 0.53
PC 0.29 0.38 0.56 0.45 37 0.085 0.155 0.55
NOTEARS ($\lambda=0.005$) 0.2 0.41 0.28 0.33 17 0.086 0.13 0.67
DAGMA ($\lambda=0.005$) 0.24 0.43 0.36 0.39 21 0.07 0.12 0.61
10000 Samples
GES 0.42 0.45 0.84 0.58 47 0.08 0.18 0.42
PC 0.26 0.32 0.6 0.42 47 0.101 0.18 0.58
NOTEARS ($\lambda=0.005$) 0.19 0.37 0.28 0.32 19 0.093 0.14 0.68
DAGMA ($\lambda=0$) 0.22 0.4 0.32 0.36 20 0.08 0.12 0.64
LLM Methods
Pairwise 0.13 0.16 0.4 0.23 61 0.165 0.215 0.76
Pairwise + 1000 samples 0.16 0.19 0.52 0.28 25 0.17 0.24 0.72
Pairwise + 10000 samples 0.14 0.16 0.56 0.25 86 0.21 0.28 0.75
Ours 0.3 0.44 0.48 0.46 27 0.07 0.13 0.54
Ours + 1000 samples 0.4 0.58 0.56 0.57 24 0.065 0.15 0.43
Ours + 10000 samples 0.46 0.59 0.68 0.63 25 0.055 0.15 0.37
Table 2: Results on the Child causal graph (20 nodes, 25 edges). The proposed method with observational statistics from 10000 samples outperforms all other methods with a F score of 0.63 and a NHD ratio of 0.37.
Method Acc. $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ F Score $\uparrow$ NPE NHD $\downarrow$ Ref. NHD NHD Ratio $\downarrow$
Numerical Methods
100 Samples
GES N/A N/A N/A N/A N/A N/A N/A N/A
PC 0.02 0.08 0.03 0.04 311 0.022 0.023 0.96
NOTEARS ($\lambda=0.01$) 0.029 0.32 0.03 0.06 76 0.1 0.11 0.94
DAGMA ($\lambda=0.01$) 0.03 0.3 0.03 0.05 77 0.10 0.11 0.95
1000 Samples
GES N/A N/A N/A N/A N/A N/A N/A N/A
PC 0.04 0.09 0.07 0.08 559 0.025 0.027 0.92
NOTEARS ($\lambda=0.01$) 0.02 0.47 0.02 0.04 36 0.34 0.35 0.96
DAGMA ($\lambda=0$) 0.03 0.098 0.040 0.057 316 0.040 0.042 0.94
10000 Samples
GES N/A N/A N/A N/A N/A N/A N/A N/A
PC N/A N/A N/A N/A N/A N/A N/A N/A
NOTEARS ($\lambda=0.005$) 0.03 0.45 0.031 0.058 53 0.19 0.20 0.94
DAGMA ($\lambda=0$) 0.033 0.14 0.040 0.063 214 0.059 0.063 0.94
LLM Methods
Pairwise N/A N/A N/A N/A N/A N/A N/A N/A
Ours 0.217 0.583 0.251 0.351 331 0.014 0.022 0.643
Table 3: Results on the Neuropathic Pain causal graph (221 nodes, 770 edges). We omit the results for GES and pairwise queries because they are intractable to use on a graph of this size. All methods except the proposed method fail catastrophically. The proposed method performs reasonably well with a F score of 0.351 and a NHD ratio of 0.643.

### v4

# Equal Contribution

@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved AI capabilities OpenAI (2023). LLMs have access to a vast amount of training data, allowing them to acquire a substantial amount of knowledge, ranging from common-sense to specific domains, such as math and science. Moreover, recent papers have argued that complex behaviors, such as writing code, generating long stories, and even some reasoning capabilities, can emerge from such large-scale training Wei et al. (2022); Rozière et al. (2023); Zhao et al. (2023); Yao et al. (2023). Impressively, LLMs can use their knowledge to generate plausible explanations given observations and even reason on counterfactuals (Wang et al., 2023; Li et al., 2023; Kıcıman et al., 2023). Given these emergent abilities to understand and learn relationships from textual corpora, a natural question arises: Can LLMs reason about causal relationships?
@@
Causal discovery involves identifying the causal relationships between different variables Pearl (2009); Peters et al. (2017). This process usually results in a Directed Acyclic Graph (DAG), with edges representing the existence and direction of causal influences. This causal graph represents the relationships between variables which govern the data generation process. The causal graph also allows us to determine how changes in one variable might influence others, and serves as the foundation for further analysis related to specific tasks, including effect inference, prediction, or attribution. Ensuring the correctness of this causal graph is essential for the reliability of any subsequent analysis.
@@
Traditional causal discovery frameworks Bello et al. (2023); Zheng et al. (2018); Meek (2023) rely on numerical observations and use statistical or continuous optimization based methods to unravel confounding and causal factors. While such rigorous numerical and abstract reasoning abilities remain challenges for LLMs (Gendron et al., 2023; Xu et al., 2023; Imani et al., 2023), recent works have shown that LLMs are capable of leveraging their internal knowledge to perform commonsense and simple logical as well as causal reasoning Wei et al. (2022); Zhao et al. (2023); Yao et al. (2023); Kıcıman et al. (2023). Inspired by these results, the proposed framework utilizes LLMs to respond to causal queries solely based on metadata, such as variable names and descriptions, without accessing the numerical observations. This is similar to the metadata and knowledge-based reasoning that is done by human domain experts when they construct causal graphs.
@@
Previous works Kıcıman et al. (2023); Choi et al. (2022); Long et al. (2023b) have shown that LLMs are capable of inferring pairwise causal relationships given only metadata. However, extending this pairwise framework to full graph discovery presents additional challenges. One straightforward approach which previous works use is to perform pairwise queries over all possible pairs of variables. However, this method is not scalable since it requires ($n\choose 2$ total queries where $n$ is the number of variables. Thus, the number of queries required grows quadratically with the number of causal variables. As a result, the experiments in these works are limited to very small causal graphs.
@@
To the best of our knowledge, we present the first LLM-based causal graph discovery method applicable to larger graphs. The proposed method is much more efficient than previous LLM-based methods, and does not require observational data (unlike numerical methods), while still achieving state-of-the-art or competitive performance on three graphs of varying sizes. In addition, we show how observational data can easily be incorporated when available to improve the proposed method’s performance.
@@ 2 Related Work
Causal graph discovery has been extensively studied in many settings Zanga & Stella (2023). Until recently, most proposed methods were statistical methods that used observational data to infer causal relationships between variables. Some examples of these are Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes & Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023). Recently, many works have proposed to use LLMs for causal graph discovery Kıcıman et al. (2023); Choi et al. (2022); Long et al. (2023b). These LLM-based methods do not rely on observational data. Instead, they utilize metadata associated with the input variables, specifically variable descriptions in natural language. This approach closely resembles how human experts construct causal graphs using their domain knowledge.
2.1 Pairwise Method
All of these works use pairwise queries to infer the causal relationship between 2 variables at a time. Given variables $A$ and $B$, there are three possible causal relationships, $A\rightarrow B$, $A\leftarrow B$, and no relationship. This query can be formulated as a question in natural language with the following prompt:
@@ 2.1 Pairwise Method
Given the above information, which of the following is the most
likely:
@@ B. Changing <B> causes a change in <A>
C. There is no causal relationship between <A> and <B>
@@ 3.1 Main Method
You are a helpful assistant to a neuropathic pain diagnosis expert.
The following factors are key variables related to neuropathic pain
diagnosis which have various causal effects on each other. Our goal
is to construct a causal graph between these variables.
@@
Now you are going to use the data to construct a causal graph. You
will start with identifying the variable(s) that are unaffected by
any other variables. Think step by step. Then, provide your final
answer (variable names only) within the tags <Answer>...</Answer>.
@@
Given <Independent Variables> is(are) not affected by any
other variable and the following causal relationships:
@@
Select the variables that are caused by <Currently Visited Node>.
Think step by step. Then, provide your final answer (variable names
only) within the tags <Answer>...</Answer>.
@@ 3.2 Adding Statistics to the Prompt
Additionally, the Pearson Correlation Coefficients between
<Currently Visited Node> and the other variables are as follows:
<A>: <Pearson Correlation Coefficient between <Currently
Visited Node> and A>
<B>: <Pearson Correlation Coefficient between <Currently
Visited Node> and B>
@@
Additionally, the Pearson Correlation Coefficient between <A> and
<B> is <Correlation Coefficient>.
@@ 4.1 Experimental Setup
We use the GPT-4 API OpenAI (2023) for all of our experiments. We evaluate the proposed approach on three causal graphs of varying sizes:
@@
    Asia Lauritzen & Spiegelhalter (1988). The Asia causal graph is a small causal graph that models lung cancer in patients having potentially recently visited Asia. The ground-truth graph has 8 nodes and 8 edges. Some examples of variables in the graph are Visit to Asia, Smoking, Bronchitis, and Lung Cancer.
@@
    Child Spiegelhalter et al. (1993). The Child causal graph is a medium-sized causal graph which models congenital heart disease in newborn babies. The ground-truth graph has 20 nodes and 25 edges. Some examples of variables in the graph are Birth Asphyxia, Lung Flow, and Chest XRay.
@@
    Neuropathic Pain Tu et al. (2019) The Neuropathic Pain causal graph is a very large causal graph consisting of 221 nodes and 770 edges. It contains relationships between different nerves and the associated symptoms that patients express. It mainly consists of symptom diagnoses, pattern diagnoses, and pathophysiological diagnoses. The symptom diagnoses describe the discomfort of patients. The pattern diagnoses identify nerve roots that lead to symptom patterns. The pathophysiological diagnoses identify discoligamentous injuries.
@@
We get the Asia and Child causal graphs and data from the bnlearn package Scutari (2010), and their descriptions from Long et al. (2023a). We get the Neuropathic Pain causal graph data and descriptions from Tu et al. (2019).
We compare the proposed method, as well as the proposed method plus observational statistics from 1000 and 10000 samples, against the numerical methods of Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes & Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023), as well as against pairwise queries, and pairwise queries plus observational statistics from 1000 and 10000 samples. We use the implementations of GES and PC from the causal-learn package Zheng et al. (2023), the implementation of NOTEARS from Zheng et al. (2018), and the implementation of DAGMA from Bello et al. (2023). For the numerical methods, we conduct experiments in the low (100 samples), medium (1000 samples), and high (10000 samples) data regimes.
@@ 4.2 Metrics
ASIA (n=8, m=8) CHILD (n=20, m=25) NEUROPATHIC (n=221, m=770)
# of Samples Method Acc. ($\uparrow$) F Score ($\uparrow$) NPE Ratio ($\downarrow$) Acc. ($\uparrow$) F Score ($\uparrow$) NPE Ratio ($\downarrow$) Acc. ($\uparrow$) F Score ($\uparrow$) NPE Ratio ($\downarrow$)
100 GES 0.23 0.38 8 0.63 0.24 0.38 22 0.62 - - - -
PC 0.33 0.5 4 0.5 0.19 0.32 19 0.68 0.02 0.04 311 0.96
NOTEARS 0.5 0.67 4 0.33 0.14 0.25 23 0.75 0.029 0.06 76 0.94
DAGMA 0.22 0.36 3 0.64 0.14 0.24 41 0.76 0.03 0.05 77 0.95
Pairwise 0.33 0.5 20 0.5 0.3 0.46 27 0.54 - - - -
Ours 0.88 0.93 7 0.067 0.3 0.46 27 0.54 0.22 0.35 331 0.64
1000 GES 0.67 0.8 7 0.2 0.31 0.44 34 0.53 - - - -
PC 0.5 0.67 7 0.33 0.29 0.45 37 0.55 0.04 0.08 559 0.92
NOTEARS 0.44 0.62 5 0.38 0.2 0.33 17 0.67 0.02 0.04 36 0.96
DAGMA 0.5 0.67 4 0.33 0.24 0.39 21 0.61 0.03 0.057 316 0.94
Pairwise 0.38 0.54 14 0.45 0.4 0.57 24 0.43 - - - -
Ours 0.8 0.89 10 0.11 0.4 0.57 24 0.43 - - - -
10000 GES 0.7 0.82 9 0.18 0.42 0.58 47 0.42 - - - -
PC 0.55 0.71 9 0.29 0.26 0.42 47 0.58 - - - -
NOTEARS 0.44 0.62 5 0.38 0.19 0.32 19 0.68 0.03 0.058 53 0.94
DAGMA 0.5 0.67 4 0.33 0.22 0.36 20 0.64 0.033 0.063 214 0.94
Pairwise 0.47 0.64 17 0.36 0.14 0.25 86 0.75 - - - -
Ours 0.8 0.89 10 0.11 0.46 0.63 25 0.37 - - - -
Table 1: Results on the Asia (8 nodes, 8 edges), Child (20 nodes, 25 edges), and Neuropathic (221 nodes, 770 edges) causal graphs. Our method obtains state-of-the-art performance on all 3 causal graphs. More results can be found in the appendix.